Consider a hypothetical economy characterized by the following behavioral equations. (5 point)
C=40-14r+0.2Y Md =30-100r
I= 24-4r MS=60 G=2000 P=2
a. Derive the equations of the IS and LM curves of the economy
1
Expert's answer
2020-06-22T11:44:12-0400

We have the information:


"C=40-14r+0.2Y \\[0.3cm]
Md =30-100r\\[0.3cm]
I= 24-4r \\[0.3cm]
MS=60 \\[0.3cm]
G=2000 \\[0.3cm]
P=2"


In the goods market:



"Y = C + I + G"

Therefore:



"Y = 40-14r+0.2Y + 24-4r + 2000 \\[0.3cm]
0.8Y = - 18r + 2064\\[0.3cm]
Y = -22.5r + 2,580...............IS"